# Strong time-periodic solutions to a multilayered fluid-structure interaction problem

### v1

@@ 4.2. The thick structural layer
$C_{\mathrm{s}}=SB_{\mathrm{s}}S^{-1}=\begin{pmatrix}-\delta^{-1}\operatorname{Id}&\operatorname{Id}\\ 0&\delta L_{0}\end{pmatrix},\enspace\text{where}\enspace\mathrm{D}(C_{\mathrm{s}})=S\mathrm{D}(B_{\mathrm{s}})=\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\times\left(\mathrm{W}^{2,q}(\Omega)^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\right).$
@@
(4.10) $\tilde{C}_{\mathrm{s}}\coloneqq\operatorname{diag}(-\delta^{-1}\operatorname{Id},\delta L_{0}),\enspace\text{with}\enspace\mathrm{D}(\tilde{C}_{\mathrm{s}})=\mathrm{D}(C_{\mathrm{s}}).$
From Subsection 4.2, we immediately infer that $-\tilde{C}_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-\tilde{C}_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$. The next goal is to transfer this property from $\tilde{C}_{\mathrm{s}}$ to $C_{\mathrm{s}}$. To this end, observe that $C_{\mathrm{s}}=\tilde{C}_{\mathrm{s}}+\tilde{P}_{\mathrm{s}}$, where
@@
In other words, $-C_{\mathrm{s}}$ is bounded with respect to a fractional power of $-\tilde{C}_{\mathrm{s}}$. It then follows from [35, Prop. 13.1] that there exists $\mu\geq 0$ such that $-C_{\mathrm{s}}+\mu=-\tilde{C}_{\mathrm{s}}-\tilde{P}_{\mathrm{s}}+\mu\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-C_{\mathrm{s}}+\mu}^{\infty}<\nicefrac{{\pi}}{{2}}$. Observe that the choice $\mu=0$ is possible, since $C_{\mathrm{s}}$ is invertible. Indeed, by $0\in\rho(L_{0})$, we have
@@
revealing that $0\in\rho(C_{\mathrm{s}})$, so $-C_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-C_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$.
Furthermore, as the $\mathcal{H}^{\infty}$-calculus is preserved under isomorphisms, see e. g. [16, Prop. 2.11(vi)], we infer that $-B_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-B_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$. Next, note that $A_{\mathrm{s}}=B_{\mathrm{s}}+P_{\mathrm{s}}$, with $P_{\mathrm{s}}=\operatorname{diag}(0,\delta^{-1}\operatorname{Id})$. Let now $(d_{1},d_{2})\in\mathrm{D}(B_{\mathrm{s}})=\mathrm{D}(A_{\mathrm{s}})$. Then we find that
@@
In other words, $-A_{\mathrm{s}}$ is a bounded perturbation of $-B_{\mathrm{s}}$, so, a priori up to a shift, the boundedness of the $\mathcal{H}^{\infty}$-calculus with angle strictly less than $\nicefrac{{\pi}}{{2}}$ transfers to $-A_{\mathrm{s}}$.
@@ 4.3. Lifting arguments
After recalling properties of the Stokes operator with homogeneous boundary conditions and the thin structural layer in Subsection 4.1, and with the understanding of the thick structure gained in Subsection 4.2, we are now in the position to tackle the coupling conditions by means of lifting arguments. Let us start with the lifting associated with the continuity of velocities at the interface, i. e., $u=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3}$ on $\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0}$. At this stage, we observe that the underlying lifting procedure has already been carried out in [41, Sec. 4.3]. However, for convenience, we recall the main steps here.
To this end, take into account the stationary Stokes problem with the corresponding boundary conditions. More precisely, noting that $\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0}=\{(s,0):s\in\omega\}$, we consider
(4.11) $-\operatorname{div}\sigma_{\mathrm{f}}(w,\psi)&=f,\enspace\operatorname{div}w=0,&&\enspace\text{in}\enspace\Omega_{\mathrm{f}},\\ w&=b,&&\enspace\text{on}\enspace\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0},\\ w&=0,&&\enspace\text{on}\enspace\Gamma_{0},\\ \int_{\Omega_{\mathrm{f}}}\psi\,\mathrm{d}y&=0.\end{aligned}\right.$
The following result is classical, see for example [57, Prop. 2.3].
[…]
Let $q\in(1,\infty)$, $f\in\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}$ and $b\in\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)$. Then there exists a unique solution $(w,\psi)\in\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\times\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})$ to (4.11).
Thanks to Subsection 4.3, we may introduce the lifting operators $D_{\mathrm{fl}}\in\mathcal{L}(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega),\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3})$ as well as $D_{\mathrm{pr}}\in\mathcal{L}(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega),\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$ defined by
(4.12) $D_{\mathrm{fl}}b\coloneqq w\enspace\text{and}\enspace D_{\mathrm{pr}}b\coloneqq\psi,$
@@
(4.13) $\Delta\varphi=0,\enspace\text{in}\enspace\Omega_{\mathrm{f}},\enspace\partial_{\nu}\varphi=c,\enspace\text{on}\enspace\partial\Omega_{\mathrm{f}},\enspace\int_{\Omega_{\mathrm{f}}}\varphi\,\mathrm{d}y=0,$
@@
$N$ $\in\mathcal{L}(\mathrm{W}_{\mathrm{m}}^{-\nicefrac{{1}}{{q}},q}(\partial\Omega_{\mathrm{f}}),\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})),\enspace\text{and}\enspace$
$N$ $\in\mathcal{L}(\mathrm{L}_{\mathrm{m}}^{q}(\partial\Omega_{\mathrm{f}}),\mathrm{W}^{1+\nicefrac{{1}}{{q}}-\varepsilon,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})),\enspace\text{for all}\enspace\varepsilon>0,$
see for example [37, Thm. 4.2 and 4.3]. In this context, we also invoke the adjusted lifting operator
(4.16) $N_{1}c_{1}=Nc,\enspace\text{where}\enspace c(y)=\left\{\begin{aligned} c_{1}(s),&\enspace\text{if}\enspace y=(s,0)\in\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0},\\ 0,&\enspace\text{if}\enspace y\in\Gamma_{0}.\end{aligned}\right.$
@@
(4.17) $\Delta\varphi=\operatorname{div}f,\enspace\text{in}\enspace\Omega_{\mathrm{f}},\enspace\partial_{\nu}\varphi=f\cdot\nu,\enspace\text{on}\enspace\partial\Omega_{\mathrm{f}},$
@@
After addressing the coupling condition on the equality of velocities, we now tackle the second kinematic coupling condition, namely the continuity of displacement, i. e., $d_{1}=P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3}$ on $\Gamma_{\mathrm{s}}$. More precisely, for $\lambda\in\mathbb{C}$, we consider the stationary problem
(4.19) $\lambda d-\mathcal{L}d&=0,&&\enspace\text{in}\enspace\Omega_{\mathrm{s}},\\ d&=b^{\prime},&&\enspace\text{on}\enspace\Gamma_{\mathrm{s}},\\ d&=0,&&\enspace\text{on}\enspace\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}}.\end{aligned}\right.$
In the lemma below, we discuss the solvability of (4.19). The discussion preceding Subsection 4.2 together with the fact that $0\in\rho(L_{0})$ by Subsection 4.2 reveals that we are in the framework of [17, Thm. 2.1]. The latter implies that for all $b^{\prime}\in\mathrm{W}^{2-\nicefrac{{1}}{{q}},q}(\omega)$, there exists a unique solution $d\in\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}$ to (4.19).
[…]
Let $q\in(1,\infty)$ and $b^{\prime}\in\mathrm{W}^{2-\nicefrac{{1}}{{q}},q}(\omega)$. Then for every $\lambda\in\mathbb{C}$ with $\operatorname{Re}\lambda\geq 0$, there exists a unique solution $d\in\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}$ to (4.19).
Subsection 4.3 allows us to define the lifting operator $D_{\mathrm{s}}\colon\mathrm{W}^{2-\nicefrac{{1}}{{q}},q}(\omega)\to\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}$ by
@@
where $d$ is the solution to (4.19) in the case $\lambda=0$. Let us observe that the aforementioned operator can also be extended to a continuous operator $\tilde{D}_{\mathrm{s}}\colon\mathrm{L}^{q}(\omega)\to\mathrm{W}^{s,q}(\Omega_{\mathrm{s}})^{3}$ for all $s\in(0,\nicefrac{{1}}{{q}})$. For the existence of a Dirichlet operator in a similar context, we also refer to [4, Sec. 7.1]. By a slight abuse of notation, we still denote the associated operator by $D_{\mathrm{s}}$ for simplicity.
@@ 4.4. The multilayered fluid-structure operator
(4.21) $T_{\mathrm{s}}(d_{1},d_{2})\coloneqq P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\right).$
@@
$u$ $\in\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}),\enspace\pi\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})),$
@@
$d_{1}$ $\in\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}),\enspace d_{2}\in\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3})\enspace\text{and}\enspace$
@@
$\mathbb{P}u^{\prime}&=A_{0}\mathbb{P}(u-D_{\mathrm{fl}}\eta_{2})+\mathbb{P}f,&&\enspace\text{in}\enspace(0,T),\\ \eta_{1}^{\prime}&=\eta_{2},&&\enspace\text{in}\enspace(0,T),\\ (\operatorname{Id}+\gamma_{\mathrm{m}}N_{1})\eta_{2}^{\prime}+P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}\eta_{1}-\Delta_{\mathrm{p}}\eta_{2}&=\gamma_{\mathrm{m}}N(\Delta\mathbb{P}u\cdot\nu)+T_{\mathrm{s}}(d_{1},d_{2})+P_{\mathrm{m}}g+\gamma_{\mathrm{m}}N_{2}f,&&\enspace\text{in}\enspace(0,T),\\ \int_{\omega}\eta_{1}(\cdot,s)\,\mathrm{d}s&=0,&&\enspace\text{in}\enspace(0,T),\\ d_{1}^{\prime}&=d_{2},&&\enspace\text{in}\enspace(0,T),\\ d_{2}^{\prime}&=L_{0}\bigl{(}d_{1}-D_{\mathrm{s}}(P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3})+\delta(d_{2}-D_{\mathrm{s}}(P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3}))\bigr{)}+h,&&\enspace\text{in}\enspace(0,T),\\ (\operatorname{Id}-\mathbb{P})u&=(\operatorname{Id}-\mathbb{P})D_{\mathrm{fl}}\eta_{2},\\ \pi&=N(\Delta\mathbb{P}u\cdot\nu)-N_{1}\eta_{2}^{\prime}+N_{2}f,\\ (\mathbb{P}u,\eta_{1},\eta_{2},d_{1},d_{2})(0)&=(\mathbb{P}u_{0},\eta_{1,0},\eta_{2,0},d_{1,0},d_{2,0}).\end{aligned}\right.$
@@
For the procedure to reformulate the Stokes equation as well as the first order system accounting for the thin structural layer, we refer to [41, Prop. 4.4]. Thus, it remains to handle the thick structural layer in the rest of the proof.
@@
[…]
@@
Before reformulating the linearized problem in operator form, we elaborate on the so-called added mass operator $M_{s}$ defined by $M_{s}\coloneqq\operatorname{Id}+\gamma_{\mathrm{m}}N_{1}$, where the lifting operator $N_{1}$ has been made precise in (4.16). For the result below, see [41, Lemma 4.5].
[…]
@@
(4.23) $\mathcal{K}(f)\coloneqq P_{\mathrm{m}}\Bigl{(}\bigl{(}\mu_{\mathrm{s}}\bigl{(}\nabla f+(\nabla f)^{\top}\bigr{)}+\lambda_{\mathrm{s}}\operatorname{div}(f)\operatorname{Id}\bigr{)}\mathrm{e}_{3}\left.\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\Bigr{)}$
@@
(4.24) $\mathrm{X}_{0}=\bigl{\{}(\mathbb{P}u,\eta_{1},\eta_{2},d_{1},d_{2})$ $\in\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\times\left(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\right)\times\mathrm{L}_{\mathrm{m}}^{q}(\omega)\times\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}\times\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3}:$

### v2

@@ 5.2. The thick structural layer
$C_{\mathrm{s}}=SB_{\mathrm{s}}S^{-1}=\begin{pmatrix}-\delta^{-1}\operatorname{Id}&\operatorname{Id}\\ 0&\delta L_{0}\end{pmatrix},\kern 5.0pt\text{where}\kern 5.0pt\mathrm{D}(C_{\mathrm{s}})=S\mathrm{D}(B_{\mathrm{s}})=\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\times\left(\mathrm{W}^{2,q}(\Omega)^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\right).$
@@
(5.12) $\tilde{C}_{\mathrm{s}}\coloneqq\operatorname{diag}(-\delta^{-1}\operatorname{Id},\delta L_{0}),\kern 5.0pt\text{with}\kern 5.0pt\mathrm{D}(\tilde{C}_{\mathrm{s}})=\mathrm{D}(C_{\mathrm{s}}).$
From Subsection 5.2, we immediately infer that $\tilde{C}_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-\tilde{C}_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$. The next goal is to transfer this property from $\tilde{C}_{\mathrm{s}}$ to $C_{\mathrm{s}}$. To this end, observe that $C_{\mathrm{s}}=\tilde{C}_{\mathrm{s}}+\tilde{P}_{\mathrm{s}}$, where
@@
In other words, $C_{\mathrm{s}}$ is bounded with respect to a fractional power of $\tilde{C}_{\mathrm{s}}$. It then follows from [42, Prop. 13.1] that there exists $\mu\geq 0$ such that $C_{\mathrm{s}}+\mu=-\tilde{C}_{\mathrm{s}}-\tilde{P}_{\mathrm{s}}+\mu\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-C_{\mathrm{s}}+\mu}^{\infty}<\nicefrac{{\pi}}{{2}}$. Observe that the choice $\mu=0$ is possible, since $C_{\mathrm{s}}$ is invertible. Indeed, by $0\in\rho(L_{0})$, we have
@@
revealing that $0\in\rho(C_{\mathrm{s}})$, so $C_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-C_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$.
Furthermore, as the $\mathcal{H}^{\infty}$-calculus is preserved under isomorphisms, see Subsection 5.1(b), we infer that $B_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with $\phi_{-B_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$. Next, note that $A_{\mathrm{s}}=B_{\mathrm{s}}+P_{\mathrm{s}}$, with $P_{\mathrm{s}}=\operatorname{diag}(0,\delta^{-1}\operatorname{Id})$. Let now $(d_{1},d_{2})\in\mathrm{D}(B_{\mathrm{s}})=\mathrm{D}(A_{\mathrm{s}})$. Then we find that
@@
In other words, $A_{\mathrm{s}}$ is a bounded perturbation of $B_{\mathrm{s}}$, so, a priori up to a shift, the boundedness of the $\mathcal{H}^{\infty}$-calculus with angle strictly less than $\nicefrac{{\pi}}{{2}}$ transfers to $A_{\mathrm{s}}$.
@@ 5.3. Lifting arguments
After recalling properties of the Stokes operator with homogeneous boundary conditions and the thin structural layer in Subsection 5.1, and with the understanding of the thick structure gained in Subsection 5.2, we are now in the position to tackle the coupling conditions by means of lifting arguments. Let us start with the lifting associated with the continuity of velocities at the interface, i.e., $u=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3}$ on $\Gamma_{0}$. At this stage, we observe that the underlying lifting procedure has already been carried out in [48, Sec. 4.3]. However, for convenience, we recall the main steps here.
To this end, take into account the stationary Stokes problem with the corresponding boundary conditions. Let us recall that $\partial\Omega_{\mathrm{f}}=\Gamma_{0}\cup\Gamma_{\mathrm{f}}$ is especially of class $\mathrm{C}^{3}$, and that $\Gamma_{0}=\{(s,0):s\in\omega\}$. We consider
(5.13) $\left\{\begin{aligned} -\operatorname{div}\sigma_{\mathrm{f}}(w,\psi)&=f,\kern 5.0pt\operatorname{div}w=0,&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\\ w&=b,&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{0},\\ w&=0,&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{\mathrm{f}},\\ \int_{\Omega_{\mathrm{f}}}\psi\,\mathrm{d}y&=0.\end{aligned}\right$.
The following result is classical, see for example [64, Prop. 2.3].
Lemma 5.10.
Let $q\in(1,\infty)$, $f\in\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}$ and $b\in\mathrm{W}_{0}^{2,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3}$. Then there exists a unique solution $(w,\psi)\in\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\times\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})$ to (5.13).
Thanks to Subsection 5.3, we may introduce the lifting operators $D_{\mathrm{fl}}\in\mathcal{L}(\mathrm{W}_{0}^{2,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3},\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3})$ as well as $D_{\mathrm{pr}}\in\mathcal{L}(\mathrm{W}_{0}^{2,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3},\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$ defined by
(5.14) $D_{\mathrm{fl}}b\coloneqq w\kern 5.0pt\text{and}\kern 5.0ptD_{\mathrm{pr}}b\coloneqq\psi,$
@@
(5.15) $\Delta\varphi=0,\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\kern 5.0pt\partial_{\nu}\varphi=c,\kern 5.0pt\text{on}\kern 5.0pt\partial\Omega_{\mathrm{f}},\kern 5.0pt\int_{\Omega_{\mathrm{f}}}\varphi\,\mathrm{d}y=0,$
@@
$N$ $\in\mathcal{L}(\mathrm{W}_{\mathrm{m}}^{-\nicefrac{{1}}{{q}},q}(\partial\Omega_{\mathrm{f}}),\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})),\kern 5.0pt\text{and}\kern 5.0pt$
$N$ $\in\mathcal{L}(\mathrm{L}_{\mathrm{m}}^{q}(\partial\Omega_{\mathrm{f}}),\mathrm{W}^{1+\nicefrac{{1}}{{q}}-\varepsilon,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})),\kern 5.0pt\text{for all}\kern 5.0pt\varepsilon>0,$
see for example [44, Thm. 4.2 and 4.3]. In this context, we also invoke the adjusted lifting operator
(5.18) $N_{1}c_{1}=Nc,\kern 5.0pt\text{where}\kern 5.0ptc(y)=\left\{\begin{aligned} c_{1}(s),&\kern 5.0pt\text{if}\kern 5.0pty=(s,0)\in\Gamma_{0},\\ 0,&\kern 5.0pt\text{if}\kern 5.0pty\in\Gamma_{\mathrm{f}}.\end{aligned}\right$.
@@
(5.19) $\Delta\varphi=\operatorname{div}f,\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\kern 5.0pt\partial_{\nu}\varphi=f\cdot\nu,\kern 5.0pt\text{on}\kern 5.0pt\partial\Omega_{\mathrm{f}},$
@@
After addressing the coupling condition on the equality of velocities, we now tackle the second kinematic coupling condition, namely the continuity of displacement, i.e., $d_{1}=P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3}$ on $\Gamma_{0}$. We recall that $\partial\Omega_{\mathrm{s}}=\Gamma_{0}\cup\Gamma_{\mathrm{s}}$ is of class $\mathrm{C}^{2}$ by assumption. For $\lambda\in\mathbb{C}$, we then consider the stationary problem
(5.21) $\left\{\begin{aligned} \lambda d-\mathcal{L}d&=0,&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{s}},\\ d&=b^{\prime},&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{0},\\ d&=0,&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{\mathrm{s}}.\end{aligned}\right$.
In the lemma below, we discuss the solvability of (5.21). The discussion preceding Subsection 5.2 together with the fact that $0\in\rho(L_{0})$ by Subsection 5.2 reveals that we are in the framework of [23, Thm. 2.1]. The latter implies that for all $b^{\prime}\in\mathrm{W}_{0}^{2,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3}$, there exists a unique solution $d\in\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}$ to (5.21).
Lemma 5.11.
Let $q\in(1,\infty)$ and $b^{\prime}\in\mathrm{W}_{0}^{2,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3}$. Then for every $\lambda\in\mathbb{C}$ with $\operatorname{Re}\lambda\geq 0$, there exists a unique solution $d\in\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}$ to (5.21).
Subsection 5.3 allows us to define the lifting operator $D_{\mathrm{s}}\colon\mathrm{W}_{0}^{2,q}(\omega)^{3}\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3}\to\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}$ by
@@
where $d$ is the solution to (5.21) in the case $\lambda=0$. Let us observe that the aforementioned operator can also be extended to a continuous operator $\tilde{D}_{\mathrm{s}}\colon\mathrm{L}_{\mathrm{m}}^{q}(\omega)^{3}\to\mathrm{W}^{s,q}(\Omega_{\mathrm{s}})^{3}$ for all $s\in(0,\nicefrac{{1}}{{q}})$. For the existence of a Dirichlet operator in a similar context, we also refer to [9, Sec. 7.1]. By a slight abuse of notation, we still denote the associated operator by $D_{\mathrm{s}}$ for simplicity.
@@ 5.4. The multilayered fluid-structure operator
(5.23) $T_{\mathrm{s}}(d_{1},d_{2})\coloneqq P_{\mathrm{m}}\left(\left.(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}_{3})\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\right).$
@@
$u$ $\in\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}),\kern 5.0pt\pi\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}})),$
@@
$d_{1}$ $\in\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}),\kern 5.0ptd_{2}\in\mathrm{W}^{1,p}(0,T;\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3})\kern 5.0pt\text{and}\kern 5.0pt$
@@
$\left\{\begin{aligned} \mathbb{P}u^{\prime}&=A_{0}\mathbb{P}(u-D_{\mathrm{fl}}\eta_{2})+\mathbb{P}f,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ \eta_{1}^{\prime}&=\eta_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ (\operatorname{Id}+\gamma_{\mathrm{m}}N_{1})\eta_{2}^{\prime}+P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}\eta_{1}-\Delta_{\mathrm{p}}\eta_{2}&=\gamma_{\mathrm{m}}N(\Delta\mathbb{P}u\cdot\nu)+T_{\mathrm{s}}(d_{1},d_{2})+P_{\mathrm{m}}g+\gamma_{\mathrm{m}}N_{2}f,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ \int_{\omega}\eta_{1}(\cdot,s)\,\mathrm{d}s&=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ d_{1}^{\prime}&=d_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ d_{2}^{\prime}&=L_{0}\bigl(d_{1}-D_{\mathrm{s}}(P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3})+\delta(d_{2}-D_{\mathrm{s}}(P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3}))\bigr)+h,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ (\operatorname{Id}-\mathbb{P})u&=(\operatorname{Id}-\mathbb{P})D_{\mathrm{fl}}\eta_{2},\\ \pi&=N(\Delta\mathbb{P}u\cdot\nu)-N_{1}\eta_{2}^{\prime}+N_{2}f,\\ (\mathbb{P}u,\eta_{1},\eta_{2},d_{1},d_{2})(0)&=(\mathbb{P}u_{0},\eta_{1,0},\eta_{2,0},d_{1,0},d_{2,0}).\end{aligned}\right$.
@@
For the procedure to reformulate the Stokes equation as well as the first order system accounting for the thin structural layer, we refer to [48, Prop. 4.4]. Thus, it remains to handle the thick structural layer in the rest of the proof.
@@
(5.24) $\left\{\begin{aligned} \partial_{t}d_{1}&=d_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{s}},\\ \partial_{t}d_{2}-\operatorname{div}\sigma_{\mathrm{s}}(d_{1},d_{2})&=h,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\Omega_{\mathrm{s}},\\ d_{1}&=b^{\prime},&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{0},\\ d_{1}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\Gamma_{\mathrm{s}},\\ d_{1}(0)&=d_{1,0},\kern 5.0ptd_{2}(0)=d_{2,0},&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{s}}.\end{aligned}\right$.
@@
Before reformulating the linearized problem in operator form, we elaborate on the so-called added mass operator $M_{s}$ defined by $M_{s}\coloneqq\operatorname{Id}+\gamma_{\mathrm{m}}N_{1}$, where the lifting operator $N_{1}$ has been made precise in (5.18). For the result below, see [48, Lemma 4.5].
Lemma 5.13.
@@
(5.25) $\mathcal{K}(f)\coloneqq P_{\mathrm{m}}\Bigl(\bigl(\mu_{\mathrm{s}}\bigl(\nabla f+(\nabla f)^{\top}\bigr)+\lambda_{\mathrm{s}}\operatorname{div}(f)\operatorname{Id}\bigr)\mathrm{e}_{3}\left.\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\Bigr)$
@@
(5.26) $\mathrm{X}_{0}=\Bigl\{(\mathbb{P}u,\eta_{1},\eta_{2},d_{1},d_{2})$ $\in\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\times\bigl(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\bigr)\times\mathrm{L}_{\mathrm{m}}^{q}(\omega)\times\mathrm{W}^{1,q}(\Omega_{\mathrm{s}})^{3}\times\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3}:$